Integrate ∫x^2+x+5/(x^2+4)(x+1)dx
1
Expert's answer
2020-03-13T09:53:08-0400

"\intop \frac{x^2+x+5}{(x^2+4)(x+1)} dx"

Let's decompose a fraction:

"\frac{x^2+x+5}{(x^2+4)(x+1)} = \frac{Kx+L}{x^2+4} + \frac{M}{x+1} = \frac{(Kx+L)(x+1)+M(x^2+4)}{(x^2+4)(x+1)} = \frac{Kx^2+Kx+Lx+L+Mx^2+4M}{(x^2+4)(x+1)} = \frac{(K+M)x^2+(K+L)x+(L+4M)}{(x^2+4)(x+1)}"

"\begin{cases}
 K+M=1 \\
K+L=1 \\
L+4M=5
\end{cases}"

"\begin{cases}
M=1-K \\
L=1-K \\
L+4M=5
\end{cases}"

"1-K+4(1-K)=5"

"1-K+4-4K=5"

"-5K=0"

"K=0"

"M=1-K=1-0=1"

"L=1-K=1-0=1"

So,

"\intop \frac{x^2+x+5}{(x^2+4)(x+1)} dx = \intop (\frac{0x+1}{x^2+4} + \frac{1}{x+1})dx = \intop (\frac{1}{x^2+4} + \frac{1}{x+1})dx = \intop \frac{dx}{x^2+4} + \intop \frac{dx}{x+1} = I_1 + I_2"

"I_1 = \intop \frac{dx}{x^2+4} = \intop \frac{dx}{x^2+2^2} = \frac{1}{2} \arctan \frac{x}{2} +C"

"I_2 = \intop \frac{dx}{x+1} = \ln (x+1) + C"

"\intop \frac{x^2+x+5}{(x^2+4)(x+1)} dx = \frac{1}{2} \arctan \frac{x}{2} + \ln (x+1) + C"